Suppose the demand for jackets is given by Qd = 120 – P, and the supply...

  1. Suppose the demand for jackets is given by Qd = 120 – P, and the supply of jackets is given by Qs = -30 + 2P.
    1. Solve for the equilibrium price.
    2. Plug the equilibrium price back into the demand equation and solve for the equilibrium quantity.
    3. Double-check your work by plugging the equilibrium price into the supply equation and solving for the equilibrium quantity. Does your answer agree with what you got in (b)?
    4. Solve for the elasticity of demand at the equilibrium point.
    5. Is demand elastic, inelastic or unit elastic? Briefly explain your answer.
Answer #1

QD = 120-P

QS = -30+2P

120-P = -30+2P

120+30 = 2P+P

150 = 3P

a) P = 150/3 = 50

b) Q = 120-50 = 70

c) When P = 50, putting it into the supply equation

Q = -30+2*50

Q = 70

Which shows that the quantity is 70 when P=50 as found in the part b

d) E = -P*P/Q

= -1*50/70

= -0.71

e) The elasticity of demand is inelastic as the value is less than one.